OpenStudy (lorenbeech):

The function f (x) has the value f (−1) = 1. The slope of the curve y = f (x) at any point is given by the expression dy/dx = (2x + 1)( y + 1) C. Use separation of variables to find an explicit or implicit formula for y = f (x), with no integrals remaining. D.Find lim f (x) x-> ∞

OpenStudy (lorenbeech):

@agent0smith

OpenStudy (agent0smith):

Separate the variables like so dy/(y+1) = (2x+1)dx Then integrate both sides. But I don't have time to stay.

OpenStudy (sshayer):

\[\frac{ dy }{ dx }=\left( 2x+1 \right)\left( y+1 \right)\] separating the variables and integrating \[\int\limits \frac{ dy }{ y+1 }= \int\limits \left( 2x+1 \right)dx+c\] \[\ln \left| y+1 \right|=\frac{ 2 x^2 }{ 2 }+x=c\] when x=-1,y=1 find c and then substitute the value of c

OpenStudy (sshayer):

by mistake typed =c,it is +c

OpenStudy (lorenbeech):

c=-x^2+1+log(2)

OpenStudy (lorenbeech):

@sshayer Is this right

OpenStudy (lorenbeech):

Or is it c=log(2)+3

OpenStudy (lorenbeech):

Could you check @agent0smith ? I think Shayer went offline

OpenStudy (agent0smith):

Looks like c=ln2

OpenStudy (lorenbeech):

Oh oops

OpenStudy (lorenbeech):

So then I'd just substitute that in and I got -2.69897

OpenStudy (agent0smith):

What? All you do it put in c into the equation for y from earlier

OpenStudy (catbuck1):

can you help me please?: The function f(x) = 4(4)x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate? f(x) = 4(4)x; growth rate 400% f(x) = 4(4)3x, growth rate 4% f(x) = 4(1.59)3x; growth rate 59% f(x) = 4(1.59)x; growth rate 4%

OpenStudy (lorenbeech):

Wait what? @agent0smith

OpenStudy (agent0smith):

The equation sshayer gave

OpenStudy (lorenbeech):

ln|y+1|=2x22+x=c this?

OpenStudy (agent0smith):

Yeah that one

OpenStudy (lorenbeech):

I did that....hm....

OpenStudy (agent0smith):

You didn't give an equation for y, you gave -2.69 or something

OpenStudy (lorenbeech):

Yikes I'm confused

OpenStudy (agent0smith):

ln|y+1|=x^2+x+c Just replace c with ln2. Then maybe solve for y.

OpenStudy (lorenbeech):

y= e^(-1+x^2) (-2-e^(1-x^2))

OpenStudy (lorenbeech):

Probably wrong again but I honestly don't know

OpenStudy (agent0smith):

I don't really have time to to check but it'd be much easier if you showed your work

OpenStudy (lorenbeech):

I plugged it into WolfFram

OpenStudy (agent0smith):

It's probably fine

OpenStudy (lorenbeech):

But it had two possible solutions

OpenStudy (lorenbeech):

Can anyone help me with d?

OpenStudy (lorenbeech):

@ShadowLegendX COuld you check this out?

satellite73 (satellite73):

i don't know much about this but maybe i could help

satellite73 (satellite73):

once you got \[\log(y+1)=x+x^2+c\] can't you solve for \(y\) via \[y+1=e^{x+x^2+c}-1=c_1e^{x+x^2}-1\]

satellite73 (satellite73):

oops i just meant \[y=c_1e^{x+x^2}-1\]

satellite73 (satellite73):

so \[f(x)=c_1e^{x+x^2}-1\\ f(-1)=c_1e^0-1=1\] making \(c_1=2\)

satellite73 (satellite73):

and so limit is infinity by your eyeballs

satellite73 (satellite73):

i am not sure the point of that last question, except that it is determined by \(c_1\)

OpenStudy (sshayer):

when x=-1 y=1 \[\ln 2=1-1+c,c=\ln 2\] ln|y+1|=x^2+x+ln2 \[\ln \left| y+1 \right|-\ln 2=x^2+x\] \[\ln \frac{ y+1 }{ 2 }=x^2+x,\frac{ y+1 }{ 2 }=e ^{x^2+x}\] \[y+1=2e ^{x^2+x}\] \[y=2 e ^{x^2+x}-1\]